OpenStudy (anonymous):

3. What is a polynomial function in standard form with zeroes 1, 2, –3, and –3? (1 point)g(x) = x4 + 3x3 –7x2 – 15x + 18 g(x) = x4 + 3x3 –7x2 + 2x + 18 g(x) = x4 – 3x3 + 7x2 + 15x + 18 g(x) = x4 – 3x3 –7x2 + 15x + 18 could you show me step by step please

OpenStudy (anonymous):

g(x) = (x - 1)(x - 2)(x + 3)(x + 3) That accounts for all of your zeros as you will see that this function is zero for all of the above zeros. Now, just multiply out.

OpenStudy (anonymous):

This fits the form of g(x) = (x - x1)(x - x2)(x - x3)(x - x4) where x1 = 1, x2 = 2, x3 = -3 = x4

OpenStudy (anonymous):

How are you doing with this?

OpenStudy (anonymous):

bad i have more like this can u show me each step then tht anser

OpenStudy (anonymous):

(x - x1)(x - x2) = x(x - x2) - x1(x - x2) = x^2 - (x)(x2) - (x)(x1) + (x1)(x2) And just keep using the distributive law of multiplication.

OpenStudy (anonymous):

i dont no that law that why i wahnt u to show each step then the ansers

OpenStudy (anonymous):

I just showed you the steps in my last post. That is the distributive law in action. Now it is up to you to try to work out the answer.